

********************************************************************
UARS WEEKLY STATUS REPORT - 28 JAN 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
***********
SUSIM
***********
SUSIM continues to gather a minimum set of daily solar scans scheduled
for periods of predicted telemetry.
***********
ACRIM
***********
Nothing receieved.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
Various options for future Narrow Band Tape Recoder operations
are being explored, and the MPG is supporting these studies.
***********
CDHF
***********
The CDHF continues to closely monitor production jobs CPU and Wallclock
usage and notify PI's of any discrepancies encountered in the processing as
a result of the UARS Spacecraft Anomaly.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
33 days WIN (5.11A) PROD (L1-L3) (520R-344C-13F=163L)
DAAC DATA:
- DAAC data created in the month of November 1999 has been loaded into the
DAAC Transfer Tables for HALOE, PEM, NMC, and UKMO.
- The CDHF will load December 1999 current day DAAC data in February 2000.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 January, 2000 (Orbit 45604, GMT Day 015) through 21
January 2000 (Orbit 45708, GMT Day 021).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 015/0103 45605 AXIS-1 On
WINDII 017/0110 45635 On
The beta angle ranged between the angles of 30.7 degrees to 59.4 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
76.0 degrees on 27 January 1999 (DOY 027).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 14 January 2000 and 21 January 2000 are provided in Appendix D.
The clock error ranged from 12.5 to -11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 06
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 09
% HALOE = 01
% PEM = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
7 Transponder A: 99/358 - 2287.496855
7 Transponder B: 99/365 - 2287.499111
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/020 show the observatory data loss to be 564 hours, 16
minutes, 10.825 seconds (an increase of 01 hours 09 minutes 20.224 seconds
since the last report period).
The increase of 01 hours 09 minutes 20.224 seconds occurred over six (5) days:
- 18 minutes 18.528 seconds on Day 014 due to an unspecified problem.
- 11 minutes 54.048 seconds on Day 015 due to an unspecified problem.
- 02 minutes 42.752 seconds on Day 016 due to an unspecified problem.
- 10 minutes 24.352 seconds on Day 019 due to an unspecified problem.
- 26 minutes 00.544 seconds on Day 020 due to an unspecified problem.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
No significant satellite operations to report.
MEETINGS
20 January, 2000 - The FOT attended a meeting to discuss the status of the tape
recorder. Three TDRSS operations were discussed in detail. The FOT was given
some action items that are to delivered by next weeks tape recorder anomaly
meeting (January 26).
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for February 15, 2000 (DOY 046). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 January 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45605 015/0103 00-014 PEM Power ON, AXIS 1
45612 015/1302 00-015 OBC Clk Adj fr 0.00 to 20.6 msec/day
45616 015/2002 00-016 NBTR NBTR NOOPS
45630 016/1824 00-017 NBTR NBTR NOOPS
45635 017/0110 00-018 WINDII Power ON
45640 017/0948 00-019 OBC Clk Adj fr 20.6 to 0.00 msec/day
45645 017/1827 00-020 NBTR NBTR NOOPS
45661 018/1912 00-021 NBTR NBTR NOOPS
45667 019/0521 00-022 HALOE Power OFF, High Temp
45667 019/0521 00-023 OBC Clk Adj fr 0.00 to 20.6 msec/day
45675 019/1837 00-024 NBTR NBTR NOOPS
45694 020/2339 00-025 NBTR NBTR Playback
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 January 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI HRDI PB CURRENT MONITOR continuous
HRPTSPBIMON Y-HI HRDI PB CURRENT MONITOR continuous
HRPTSPBIMON Y-HI HRDI PB CURRENT MONITOR 45634-41
PWICDH Y-LO CDH POWER LOW 45655
PEAXTHERM6E Y-HI PEM AXIS THERMISTOR 6 45704,05
PWICDH Y-LO CDH CURRENT 45704
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 January 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45612 015/13:08 (15 Jan 00) from 00.0 to 20.6 msec/day
45640 017/09:50 (17 Jan 00) from 20.6 to 00.0 msec/day
45667 019/05:23 (19 Jan 00) from 20.6 to 00.0 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 January 2000
14 Jan 2000 - GMT Day 014 Beta = 30.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------- -----
1 -50.4/-78.4 -3.05/-4.31 +0.0/+0.0 16.2 0.000 0.0
2 +11.2/-78.4 +6.44/+4.83 +30.4/-15.2 24.2 1.023 22.5
3 +33.6/-67.2 +5.15/+3.87 +29.6/-16.0 24.2 1.017 22.3
21 Jan 2000 - GMT Day 021 Beta = 59.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ----- ------ ------
1 -61.6/-89.6 -3.05/-4.94 +0.0/+0.0 16.2 0.000 0.0
2 +11.2/-22.4 +4.51/+2.60 +24.0/-15.2 26.4 1.040 17.6
3 +33.6/-11.2 +3.23/+1.66 +23.2/-14.4 26.4 1.038 17.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ----- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 January 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-001 FOT 45702 021/1325 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
00-001 FOT See list January 14170 Generic Late Acquisition
SEVNTFW00/03